Marketable Debt Securities Use the financial statement effects template to record the accounts and amounts for the following four transactions involving investments in marketable debt securities classified as available-for-sale securities. Assume that these transactions occur in 2016 (before the new rules for securities went into effect). a. Loudder Inc. purchases 4,000 bonds with a face value of $1,000 per bond. The bonds are purchased at par for cash and pay interest at a semi-annual rate of 4%. b. Loudder receives semi-annual cash interest of $160,000. c. Year-end fair value of the bonds is $978 per bond. d. Shortly after year-end, Loudder sells all 4,000 bonds for $970 per bond. Use negative signs with answers, if appropriate. Balance Sheet Income Statement Noncash Assets Contrib. Captial Cat Earned Capital Capelet Transaction Cash Asset + = Liabilities + + Revenues . Expenses = Net income Loudder purchases bonds. < < 160,000 160,000 160,000 o 160,000 Loudder receives cash interest. < < xoxo o xo «• « • < • < o Bonds year-end fair value is determined. . < < O XO < o o O Loudder sells all bonds xo Check
Answer #1

a. Loudder purchases bons, then in cash asset column -4000000, the reason is that cash out flow. and Non cash Asset will increase by 4000000.

b. Loudder received cash interest of 4% cash will icrease by 160000, ans retained profit or earned capital will increase by 160000 in Balance sheet. Revenue will increase by same amount and net income also will increase by 160000.

c. Bond's year-end fair value is determined:- The amount will be (1000-978) X 4000 = 88000. The Non cash Asset will decrease by 88000. so show as -88000. Capital earned will also decrease by 88000, so show as -88000. And will not affect given columns of income statements. Because this item will show in other comprehensive income under US GAAP after net income.

d. Loudder sells all these bonds at $978 per bond:- The Non cash Asset will be decreased by 3912000 and show as -3912000, and Cash asset will increase by 978 X 4000 = 3880000. Capital earned will decrease by (978 - 970) X 4000 =32000 and show as -32000, Realised loss willbe (978 - 970) X 4000 =32000? and show in Revenue column as -32000. And Net Income also will decrease by same amount of 32000 and show as -32000.

Please rate the answer maximum if you get the answer and satisfied. If you remains any doubts on this answer, please leave a comment and it will be cleared.
Thank you,…